OpenStudy (anonymous):

Two angles of a triangle measure 58 and 104. What is the measure of the third angle? A. 18 B. 28 C. 32 D. 76

OpenStudy (anonymous):

A. 18

OpenStudy (arabpride):

58 + 104 + x = 180 162 + x = 180 -162 -162 x = 18

OpenStudy (anonymous):

OK thank you everyone

OpenStudy (arabpride):

Welcome! ANy Time *^.^*

OpenStudy (anonymous):

A triangle is 180 degrees. Subtract 104 from 180, that leaves you with 76, then subtract 58 from 76, and that leaves you with 18.